Suppose the Home firm is considering whether to enter the Foreign market. Assume that the Home firm has the following costs and demand:

Fixed costs = $100 Marginal costs = $15 per unit Local price = $30 Local quantity = 40 Export price = $25 Export quantity = 20

  1. Calculate the firm’s total costs from selling only in the local market
  2. What is the firm’s average cost from selling only in the local market?
  3. Calculate the firm’s profit from selling only in the local market.
  4. d. Should the Home firm enter the Foreign market? Briefly explain why.
  5. e. Calculate the firm’s profit from selling to both markets.
  6. f. Is the Home firm dumping? Briefly explain.
Answer #1

(a)

Total cost ($) = FC + Q x MC = 100 + 15 x 40 = 100 + 600 = 700

(b)

Average cost ($) = Total cost / Output = 700/40 = 17.5

(c)

Revenue ($) = P x Q = 30 x 40 = 1200

Profit ($) = Revenue - Total cost = 1200 - 700 = 500

(d)

Export price is higher than average cost in local market. So firm should enter foreign market.

(e)

Export market revenue ($) = 25 x 20 = 500

Aggregate revenue ($) = 1200 + 500 = 1700

Aggregate cost ($) = Local cost + Export cost = 700 + (15 x 20) = 700 + 300 = 1000

Aggregate profit ($) = 1700 - 1000 = 700

(f)

Since export price is higher than local marginal cost and local average cost, the firm is not dumping.